If one integer is 3 more than another integer, the differencve in their squares is 6 more than 5 times the larger integer. What are the integers.
I think I am on the wrong tract, please help: 5(3x^2+6)=y^2
Thank you

Answer by edjones(8007)

Let the integers be x and x+3
(x+3)^2-x^2=5(x+3)+6
x^2+6x+9-x^2=5x+15+6
6x+9=5x+21
x=12
.
Ed